OpenStudy (matimaticas):

Find the length of the arc when: r = 55 inches, theta = 50 degrees

OpenStudy (anonymous):

\[\theta=\frac{ l }{ r },\theta ~is ~\in~radians.\]

OpenStudy (matimaticas):

@surjithayer what is the l in that formula?

OpenStudy (anonymous):

\[180~ degree=\pi~ radians. \] \[l=r \theta ,l=length~ of~ arc\]

OpenStudy (matimaticas):

well the answer choices are A. 46.15 inches B. 49.29 inches C. 48 inches D. 50.43 inches

OpenStudy (noelgreco):

Until you memorize formulas, the easiest way to think of arc length is to consider the fraction of the circumference with which you're dealing. \[2 \pi r=C\] Fraction of circumference is 50/360 degrees. Multiply that by the circumference for arc length.

OpenStudy (noelgreco):

Remember, you only have two significant figures, so you have to round

OpenStudy (jhannybean):

To convert from degrees to radians : \(50 \ \cdot \frac{\pi}{180}\) Just saying :P

OpenStudy (jhannybean):

If that's useful in anyway for you.

OpenStudy (matimaticas):

I got 48 :) thanks... all of you

OpenStudy (anonymous):

can someone help me

OpenStudy (jhannybean):

Put up your own question in your own thread :)

OpenStudy (anonymous):

okay sorry just my first time don't know how to use this

OpenStudy (jhannybean):

that's ok! Don't worry :)